OpenStudy (anonymous):

TELL ME THE ANSWER FOR A MEDAL!! If g(x) = x2 + 2, find g(3). 9 8 11 6

OpenStudy (anonymous):

asap please?!

OpenStudy (anonymous):

Its 8

OpenStudy (anonymous):

can you explain how you got your answer please?

OpenStudy (anonymous):

g(3)=(3)2+2 g(3)=6+2 g(3)=8

OpenStudy (phi):

@Danielle_Ello Please read the Code of Conduct. You will be banned from this site if you do not follow the Code of Conduct.

OpenStudy (anonymous):

@prettyunicorn okay thanks:) @phi sorry what's the 'Code of Conduct'?

OpenStudy (phi):

The link is in the lower left below the posted questions.

OpenStudy (anonymous):

k, gotcha:)